The sum of 2 consecutive even integers is n. What is the sum of the next two consecutive integers greater than n?: The sum of 2 consecutive even integers is n. What is the sum of the next two consecutive integers greater than n?
Answer by stanbon(18754)

The sum of 2 consecutive even integers is n. What is the sum of the next two consecutive integers greater than n?
------------------------------
1st: 2x
2nd: 2x+2
------------------------
2x + 2x+2 = n
-----------------
3th: 2x+4
4th: 2x+6
-----------------
Sum = 4x+10 = n + 4
==========================
Cheers,
Stan H.
=========================================